OpenStudy (anonymous):

five years ago, ross was n times as old as amanda was. if amanda is now 19 years old, how old is ross now in terms of n A) 14N-5 B)14N+5 c)19N+5 D)15N+5 E)19N-5

OpenStudy (anonymous):

let ross age now be r and amanda's a=19 then before 5 years: r-5=n*(a-5) then we have r-5=n*14 r=14n+5 there you have it